An X-ray line of wavelength 0.53832A0 is found to be emitted from an X− ray tube with (Z= 30) target in addition to the characteristic Kα line of Zinc of wavelength 1.43603A0 . If the unkown line is due to an impurity in the target, obtain the atomic number of the impurity.


1
Expert's answer
2021-12-21T12:24:14-0500

"\frac{\lambda_1}{\lambda_2}=\frac{z_2^2}{z_1^2},"

"z_2=z_1\sqrt{\frac{\lambda_1}{\lambda_2}}=18."